# VideoDrafter: Content-Consistent Multi-Scene Video Generation with LLM

### v1

@@ 4.3 Implementation Details
We implement our VideoDrafter using the Diffusers codebase on the PyTorch platform.
[…]
Training stage of VideoDrafter-Vid. VideoDrafter-Vid model is developed based on the Stable Diffusion XL framework by inserting temporal attentions and temporal convolutions. The training is carried out on the WebVid-10M and HD-VG-130M datasets. For each video, we randomly sample a 16-frame clip with the resolution of 320$\times$512 and an FPS of 8. The middle frame of the clip is utilized as the scene-reference image. Each minibatch consists of 128 video clips implemented on 64 A100 GPUs in parallel. We utilize the AdamW optimizer with a fixed learning rate of $3\times 10^{-6}$ for 480K iterations.
Table 1: Performance comparisons of VideoDrafter-Img variants with different input references on MSR-VTT validation set.
@@
4.4 Experimental Analysis of VideoDrafter
Evaluation on VideoDrafter-Img. We first verify the effectiveness of VideoDrafter-Img in aligning with the input entity reference images. To this end, we take the prompts from MSR-VTT validation set. The input foreground and background reference images are produced by using LLM and Stable Diffusion model. We validate the generated images on the measure of foreground similarity (FG-SIM) and background similarity (BG-SIM), which are the CLIP-similarity values with the foreground and background reference images, respectively. Table 1 lists the performance comparisons of different VideoDrafter variants by leveraging different input references. Overall, the use of foreground/background reference image as guidance leads to higher FG-SIM/BG-SIM values. The combination of both reference images achieves the highest FG-SIM of 0.8102 and BG-SIM of 0.5861. It is worth noting that all variants obtain close CLIPSIM values, indicating that the involvement of reference images does not affect the alignment with the input prompts. Figure 5 showcases six generated images by different VideoDrafter-Img variants with various reference images. The results demonstrate the advantage of VideoDrafter-Img to align with the entity reference images.
Figure 5: The examples of the foreground and background reference images and the generated scene-reference image by the VideoDrafter-Img variants.
@@ 4.4 Experimental Analysis of VideoDrafter
${}^{-}$ 157.3 98.5
RF+VideoDrafter-Vid 116.5 98.8
Figure 6: Examples of generated multi-scene videos by ModelScopeT2V [51], VideoDirectorGPT [22] and our VideoDrafter utilizing a multi-scene prompt from the Coref-SV dataset. For each video, only the first four scenes are given. The results of VideoDirectorGPT are provided in the project webpage and thus with bounding box annotation.
Evaluation on VideoDrafter-Vid. Next, we assess the visual quality of the single-scene videos generated by VideoDrafter-Vid. We exploit the real frame from the WebVid-10M validation set as the scene-reference image irrespective of the generation quality, and produce a video using the corresponding text prompt, which is referred to as RF+VideoDrafter-Vid. We compare our proposal with two baseline models of RF+VideoComposer and ${}^{-}$. The former employs a pre-trained image animation model by VideoComposer [53] to convert the real frame into a video clip. The latter disables the action guidance in VideoDrafter-Vid. Table 2 presents the performance comparisons for single-scene video generation on the WebVid-10M dataset. With the same scene-reference images, ${}^{-}$ outperforms VideoComposer relatively by 31.9% FVD and 2.7% frame consistency. This improvement is attributed to the deep network architecture and large-scale training set. The performance is further enhanced to 116.5 FVD and 98.8 frame consistency by RF+VideoDrafter-Vid, verifying the superiority of involving action category guidance to improve visual quality and intra-scene consistency.
@@
SD+VideoDrafter-Vid 11.9 381
RF+VideoComposer [53] 31.3 208
RF+VideoDrafter-Vid 10.8 133
Similar performance trends are observed on MSR-VTT dataset, as summarized in Table 3. The methods in this table are grouped into two categories: the methods with or without real frame (RF) as reference. To compare with the generation models without RF, we develop a two-step solution that first generates the scene-reference image by Stable Diffusion, and then converts the image into a video clip by VideoDrafter-Vid, which is denoted as SD+VideoDrafter-Vid. Specifically, VideoDrafter-Vid attains the best FVD on both settings with and without a real frame as reference. SD+VideoDrafter-Vid is slightly inferior to ModelScopeT2V in FID. We speculate that this may be the result of not optimizing Stable Diffusion on video frames, resulting in poorer frame quality against ModelScopeT2V. Nevertheless, SD+VideoDrafter-Vid apparently surpasses ModelScopeT2V in FVD, validating the video-level quality by VideoDrafter-Vid.
Table 4: Performance comparisons for multi-scene video generation on ActivityNet Caption dataset.
@@
VideoDrafter w/o Ref. 17.3 624 50.8
VideoDrafter 13.2 395 75.1
Table 5: Performance comparisons for multi-scene video generation on Coref-SV dataset.
@@
VideoDrafter w/o Ref. 0.3103 40.9
VideoDrafter 0.3304 77.3
Figure 7: Examples of generated multi-scene videos by VideoDrafter on MSR-VTT. For each video, only the first four scenes are given.
@@ 4.5 Evaluations on Multi-Scene Video Generation
We validate VideoDrafter for multi-scene video generation on ActivityNet Captions and Coref-SV datasets. Both of the datasets consist of multi-scene prompts, which necessitate the LLM to write the video script based on the given prompt of each scene. We compare with three approaches: ModelScopeT2V, VideoDirectorGPT and VideoDrafter w/o Ref. by disabling the reference images in VideoDrafter. Table 4 details the performance comparisons on ActivityNet Captions. As indicated by the results in the table, VideoDrafter exhibits superior visual quality and better cross-scene consistency. Specifically, VideoDrafter surpasses VideoDrafter w/o Ref. by 24.3 scene consistency, which essentially verifies the effectiveness of incorporating entity reference images. Moreover, VideoDrafter leads to 10.3 and 29.1 improvements in scene consistency over VideoDirectorGPT and ModelScopeT2V, respectively. Similar results are also observed on Coref-SV dataset, as summarized in Table 5. Note that as Coref-SV only offers prompts without the corresponding videos, FID and FVD cannot be measured for this case. As shown in the table, VideoDrafter again achieves the highest cross-scene consistency of 77.3, making an absolute improvement of 39.4 and 34.5 over ModelScopeT2V and VideoDirectorGPT. Figure 6 showcases an example of generated four-scene videos by different approaches on Coref-SV, manifesting the ability of VideoDrafter on generating visually similar entities (e.g., mouse/garden) across scenes. Figure 8 further shows an example of multi-scene video generation by VideoDrafter using the real images as entity reference images, which demonstrates the potential of VideoDrafter in customizing the generated objects or environments.
Figure 8: One example of generated multi-scene videos by our VideoDrafter using the real images as entity reference images.
4.6 Human Evaluation
In this section, we conduct a human study to evaluate the entire process of generating multi-scene video from a single prompt. We compare our VideoDrafter with four approaches: ModelScopeT2V w/o LLM and VideoDrafter w/o Ref. w/o LLM to generate five scenes by duplicating the input prompt, ModelScopeT2V w/ LLM and VideoDrafter w/o Ref. to utilize LLM to provide video script as described in Sec. 3.1 while generate each scene individually. We invite 12 evaluators and randomly select 100 prompts from MSR-VTT validation set for human evaluation. We show all the evaluators the five videos generated by each approach plus the given prompt and ask them to rank the five videos from 1 to 5 (good to bad) with respect to the three criteria: visual quality (VQ), logical coherence (LC) and content consistency (CC). For each approach, we average the ranking on each criterion of all the generated videos. As indicated by the results in Table 6, the study proves the impact of LLM generated video script and entity reference images to improve logical coherence and content consistency, respectively. Figure 7 illustrates the examples of the generated multi-scene videos by our VideoDrafter.
@@ 4.6 Human Evaluation
VideoDrafter w/o Ref. w/o LLM 2.0 3.0 2.3
VideoDrafter w/o Ref. 2.4 2.3 3.4
VideoDrafter 1.6 1.2 1.2
@@ 5 Conclusions
We have presented a new VideoDrafter framework for content-consistent multi-scene video generation. VideoDrafter involves LLM to benefit from the logical knowledge learnt behind and rewrite the input prompt into a multi-scene video script. Then, VideoDrafter identifies common entities throughout the script and generates a reference image for each entity, which serves as the link across scenes to ensure the appearance consistency. To produce a multi-scene video, VideoDrafter devises two diffusion models of VideoDrafter-Img and VideoDrafter-Vid. VideoDrafter-Img creates a scene-reference image for each scene based on the corresponding event prompt and entity reference images. VideoDrafter-Vid converts the scene-reference image into a video clip conditioning on the specific action and camera movement. Extensive evaluations on four video benchmarks demonstrate the superior visual quality and content consistency by VideoDrafter over SOTA models.

### v2

@@ 4.3 Implementation Details
We implement the proposed VideoStudio using the Diffusers codebase on the platform of PyTorch.
Training stage of VideoStudio-Img. VideoStudio-Img is originated from the Stable Diffusion v2.1 model by incorporating two additional cross-attention modules. These modules are initialized from scratch and trained on the text-image pairs from LAION-2B dataset, while other parts of the network are frozen. For each image, we randomly sample a 512$\times$512 patch cropped from the original image, and utilize the U2-Net model to segment the foreground area of each patch. The isolated foreground and background areas serve as the foreground and background reference images, respectively, to guide the generation of the input patch. We set each minibatch as 512 patches that are processed on 64 A100 GPUs in parallel. The parameters of the model are optimized by AdamW optimizer with a fixed learning rate of $1\times 10^{-4}$ for 20K iterations.
Training stage of VideoStudio-Vid. VideoStudio-Vid is developed based on the Stable Diffusion XL architecture by inserting temporal attentions and temporal convolutions. The training is carried out on the WebVid-10M and HD-VG-130M datasets. For each video, we randomly sample a 16-frame clip with the resolution of 320$\times$512 and an FPS of 8. The middle frame of the clip is utilized as the scene-reference image. Each minibatch consists of 128 video clips implemented on 64 A100 GPUs in parallel. We utilize the AdamW optimizer with a fixed learning rate of $3\times 10^{-6}$ for 480K iterations.
4.4 Experimental Analysis of VideoStudio
Evaluation on VideoStudio-Img. We first verify the efficacy of VideoStudio-Img in aligning with the input entity reference images. To this end, we take the prompts from MSR-VTT validation set. The input foreground and background reference images are produced by using LLM and Stable Diffusion model. We validate the generated images on the measure of foreground similarity (FG-SIM) and background similarity (BG-SIM), which are the CLIP-similarity values with the foreground and background reference images, respectively. Table 5 lists the performance comparisons of IP-Adapter [64] and different VideoStudio-Img variants by leveraging different input references.
Figure 4: Performance comparisons of IP-Adapter [64] and VideoStudio-Img variants with different input references on the MSR-VTT validation set.
@@ 4.4 Experimental Analysis of VideoStudio
Figure 5: Examples of the foreground and background reference images and the generated scene-reference image by the VideoStudio-Img variants.
@@
               RF+DynamiCrafter [60]                176.8                97.5
               RF+SVD [2]                153.0                98.7
               RF+VideoStudio-Vid-                157.3                98.5
               RF+VideoStudio-Vid                116.5                98.8
Specifically, the use of foreground/background reference image as guidance leads to higher FG-SIM/BG-SIM values comparing to IP-Adapter or not leveraging reference images. Though both of IP-Adapter and VideoStudio-Img exploit additional cross-attention to maintain visual contents in image diffusion, our VideoStudio-Img is devised for a more complex scenario to specify foreground objects and background. There are two major differences: 1) We pre-segment the foreground/background of the reference images to avoid the visual content interference; 2) IP-Adapter extracts global image features from CLIP, while ours utilizes local image tokens from CLIP to improve spatial discrimination in local regions. As indicated by the results, emphasizing the feature learning of local region on the more clean (masked) foreground/background reference image does benefit the visual alignment. Furthermore, the combination of both reference images achieves the highest FG-SIM of $0.8102$ and BG-SIM of $0.5861$. Figure 5 showcases four generated images by different VideoStudio-Img variants with various reference images. The results demonstrate the advantage of VideoStudio-Img to align with the visual contents in the entity reference images.
Evaluation on VideoStudio-Vid. Next, we assess the visual quality of the single-scene videos generated by VideoStudio-Vid. We exploit the real frame from the WebVid-10M validation set as the scene-reference image irrespective of the generation quality, and produce a video using the corresponding text prompt, which is referred to as RF+VideoStudio-Vid. We compare our proposal with five image-to-video diffusion models and one variant of VideoStudio-Vid, i.e., RF+VideoStudio-Vid-, which disables the action guidance in VideoStudio-Vid. Table 1 presents the performance comparisons for single-scene video generation on the WebVid-10M dataset. With the same scene-reference images, VideoStudio-Vid- outperforms most image-to-video approaches and obtains comparable FVD performance with the strong baseline SVD. The competitive result is attributed to the deep network architecture and large-scale training set. The performance is further enhanced to $116.5$ FVD and $98.8$ frame consistency by RF+VideoStudio-Vid, verifying the superiority of involving action category guidance to improve visual quality and intra-scene consistency.
@@
               SD+VideoStudio-Vid                11.9                381
               RF+VideoCrafter [5]                ✓                45.0                339
               RF+I2VGen-XL [69]                ✓                37.4                264
               RF+VideoComposer [57]                ✓                31.3                208
               RF+DynamiCrafter [60]                ✓                26.1                196
               RF+SVD [2]                ✓                15.3                172
               RF+VideoStudio-Vid                ✓                10.8                133
Figure 6: Examples of generated multi-scene videos by ModelScopeT2V [55], VideoDirectorGPT [26] and our VideoStudio utilizing a multi-scene prompt from the Coref-SV dataset. For each video, only the first four scenes are given. The results of VideoDirectorGPT are provided in the project webpage and thus with bounding box annotation.
Similar performance trends are observed on MSR-VTT dataset, as summarized in Table 2. The methods in this table are grouped into two categories: the methods with or without real frame (RF) as reference. To compare with the generation models without RF, we develop a two-step solution that first generates the scene-reference image by Stable Diffusion, and then converts the image into a video clip by VideoStudio-Vid, which is denoted as SD+VideoStudio-Vid. Specifically, VideoStudio-Vid attains the best FVD on both settings with and without a real frame as reference. SD+VideoStudio-Vid is slightly inferior to ModelScopeT2V in FID. We speculate that this may be the result of not optimizing Stable Diffusion on video frames, resulting in poorer frame quality against ModelScopeT2V. Nevertheless, SD+VideoStudio-Vid apparently surpasses ModelScopeT2V in FVD, validating the video-level quality by VideoStudio-Vid.
To evaluate the effectiveness of the action category condition for motion generation, we additionally implement an ablation study on the recent released VBench [19] benchmark. We measure the action score in VBench to assess whether human subjects can accurately execute the specific action mentioned in the text prompts. By using the action category as the condition in video diffusion, the action score of VideoStudio-Vid is improved from $90.3\%$ to $96.5\%$, indicating the efficacy of action category condition to emphasize motion patterns.
Table 3: Performance comparisons for multi-scene video generation on ActivityNet Captions dataset.
@@
            VideoStudio w/o Ref.             17.3             624             50.8
            VideoStudio             13.2             395             75.1
Figure 7: Two examples of generated multi-scene videos by our VideoStudio using the real images as entity reference images.
Table 4: Performance comparisons for multi-scene video generation on Coref-SV.
@@
            VideoStudio w/o Ref.             0.3103             40.9
            VideoStudio             0.3304             77.3
@@ 4.5 Evaluations on Multi-Scene Video Generation
We validate VideoStudio for multi-scene video generation on ActivityNet Captions and Coref-SV datasets. Both of the datasets consist of multi-scene prompts, which necessitate the LLM to write the video script based on the given prompt of each scene. We compare with three approaches: ModelScopeT2V, VideoDirectorGPT and VideoStudio w/o Ref. by disabling the reference images in VideoStudio. Table 3 details the performance comparisons on ActivityNet Captions. As indicated by the results in the table, VideoStudio exhibits superior visual quality and better cross-scene consistency. Specifically, VideoStudio surpasses VideoStudio w/o Ref. by $24.3$ scene consistency, which essentially verifies the effectiveness of incorporating entity reference images. Moreover, VideoStudio leads to $10.3$ and $29.1$ improvements in scene consistency over VideoDirectorGPT and ModelScopeT2V, respectively. Similar results are also observed on Coref-SV dataset, as summarized in Table 4. Note that as Coref-SV only offers prompts without the corresponding videos, FID and FVD cannot be measured for this case. As shown in the table, VideoStudio again achieves the highest cross-scene consistency of $77.3$, making an absolute improvement of $39.4$ and $34.5$ over ModelScopeT2V and VideoDirectorGPT. Figure 6 showcases an example of generated four-scene videos by different approaches on Coref-SV, manifesting the ability of VideoStudio on generating visually similar entities (e.g., mouse/garden) across scenes. Figure 7 further shows two examples of multi-scene video generation by VideoStudio using the real images as entity reference images, which demonstrates the potential of VideoStudio in customizing the generated objects or environments.
Figure 8: Examples of generated multi-scene videos by VideoStudio on MSR-VTT. For each video, only the first four scenes are given.
@@
            VideoStudio w/o Ref. w/o LLM             2.0             3.0             2.3
            VideoStudio w/o Ref.             2.4             2.3             3.4
            VideoStudio             1.6             1.2             1.2
4.6 Human Evaluation
Multi-Scene Video Quality. In this section, we conduct a human study to evaluate the entire process of generating multi-scene video from a single prompt. We compare our VideoStudio with four approaches: ModelScopeT2V w/o LLM and VideoStudio w/o Ref. w/o LLM to generate five scenes by duplicating the input prompt, ModelScopeT2V w/ LLM and VideoStudio w/o Ref. to utilize LLM to provide video script as described in Sec. 3.1 while generate each scene individually. We invite 12 evaluators and randomly select $100$ prompts from MSR-VTT validation set for human evaluation. We show all the evaluators the five videos generated by each approach plus the given prompt and ask them to rank the five videos from 1 to 5 (good to bad) with respect to the three criteria: visual quality (VQ), logical coherence (LC) and content consistency (CC). For each approach, we average the ranking on each criterion of all the generated videos. As indicated by the results in Table 5, the study proves the impact of LLM in generating video script and entity reference images to improve logical coherence and content consistency, respectively. Figure 8 illustrates the examples of the generated multi-scene videos by our VideoStudio.
Table 6: User preferences on script/videos by using different LLMs in VideoStudio.
ChatGLM3-6B [8] GPT-4 [38] Tie
                      Video Script 25% 37% 38%
                      Multi-Scene Video 20% 21% 59%
Different LLMs. To further investigate the effectiveness of different LLMs for multi-scene video generation, we carried out an ablation study on a variant of VideoStudio with GPT-4 [38] in Table 6. Evaluators vote on the preferring video text script by using ChatGLM3-6B and GPT-4, and the corresponding multi-scene videos generated by VideoStudio. “Tie” refers to a close preference. The results indicate that the video script generated by GPT-4 is of higher quality than ChatGLM3-6B. This is not surprising given the significantly larger parameters of GPT-4 ($\sim$1T v.s. 6B). Nevertheless, the voting on multi-scene videos is comparable, showing that the use of an open-source LLM does not affect video quality much. Our exploitation of open-source LLM leads to an elegant view of how responses of light-weight LLM could be improved for video script generation.
@@ 5 Conclusions
We have presented a new VideoStudio framework for consistent-content and multi-scene video generation. VideoStudio involves LLM to benefit from the logical knowledge learnt behind and rewrite the input prompt into a multi-scene video script. Then, VideoStudio identifies common entities throughout the script and generates a reference image for each entity, which serves as the link across scenes to ensure the appearance consistency. To produce a multi-scene video, VideoStudio devises two diffusion models of VideoStudio-Img and VideoStudio-Vid. VideoStudio-Img creates a scene-reference image for each scene based on the corresponding event prompt and entity reference images. VideoStudio-Vid converts the scene-reference image into a video clip conditioning on the specific action and camera movement. Extensive evaluations on four video benchmarks demonstrate the superior visual quality and content consistency by VideoStudio over SOTA models.